# https://www.usc.edu/

### v1

@@ 1 Introduction
The §1 presents a brief introduction that this paper aims to resolve i.e. to delineate a classification for NeRF models and discuss about its history and future. The remainder of this paper has been organized as follow. §2 discusses the fundamental and basic concepts and principles of interest in relation to the research related to Neural Rendering and Novel View Synthesis. §3 discusses the various neural architectures before the conceptions of NeRFs. It further discusses various surveys that have been put forward in this domain and the need for this survey. §4 discusses the chronological description of the models, extensions, and improvements over Neural Radiance Fields. §5 briefly discusses the datasets that have been leveraged till now in the field of NeRFs. Further, these are also then taxonomized according to different objectives (discussed in §6), different evaluation metrics discussed in §7), and different applications (discussed in §8). All sections from §58 discuss the research in the past couple of years in the field of NeRFs (limited till December 2022). Finally, the future prospects of this family of models is explored in §9. §10 concludes the paper and discusses some future prospects for this research.
@@ 2.1 Computer Graphics
[…]
    where $v||$ denotes the magnitude of vector $v$. The resulting $\mathcal{N}$ vector is a unit vector that points in the direction of the surface normal.
@@
[…]
@@
    However, when the pitch angle approaches $\pm 90\degree$, the rotation around the y-axis (yaw) and the rotation around the z-axis (roll) become aligned, resulting in a loss of a degree of freedom. This is gimbal lock, and it can cause unexpected or incorrect behavior in animations and simulations.
@@ 2.2 Mathematics and Geometry
[…]
      where $p$ is 3D point, $\textrm{{MLP}}(p)$ is MLP output applied to $p$, and $f(p)||_{2}$ is the L2 norm of the MLP output. The NSDF output is a unit vector that points towards the nearest surface of the object. There are different types of NSDF used in computer vision and robotics applications. Some of them are:
@@
[…]
    where $x$ is 3D vector for point in space, $x||$ is its Euclidean distance, and $\sigma$ scales and controls the size of Gaussiann function.
@@
  • Discrete Fourier Transform is a mathematical technique that is a discrete variant of the Fourier Transform discussed above. It transforms a discrete-time sequence of N complex numbers $x[n]$ into a sequence of N complex numbers $X[k]$ representing the frequency content of the signal. It is defined by the eqn 47 where $k$ is frequency index ranging from 0 to N1N-1.
@@
    It can be calculated using matrix multiplication of the $N\times N$ DFT matrix $F$ and the N1N-1 column vectors.
@@
    The equation for positional encoding is given in eqn 48 where $\text{PE}{(pos,2i)}$ and $\text{PE}{(pos,2i+1)}$ are the $2i$-th and $(2i+1)$-th elements of the positional encoding vector for the pospos-th position in the sequence, $d_{\text{model}}$ is the dimensionality of the model, and $i$ is the index of the embedding dimension.
@@ 3 History of Neural Rendering and NeRFs
LLFF (Local Light Field Fusion) (example depicted in fig. 4(a)) is one prior work done by Mildenhall et al [147], some of whose authors were also contributors for the work under focus in this survey. LLFFs leverage MPI (Multiplane Image) scene representations [269] for view synthesis from an irregular grid of sampled views for expanding each sampled view into a local light fields [122] and then renders novel views by fusing these fields. LLFFs achieve perceptual quality similar to Nyquist Rate Sampling (with $4000\times$ less views) by creating a bound for real world scenes creation/rendering which is specifies the density of sample views for a given scene by extending the traditional plenoptic sampling [22] theory. Here, the LLFFs result in a quadratic decrease in the number of input views based on the number of predicted planes for each layered scene representation and still maintain the Nyquist level performance by specializing to the subset of natural scenes. Further, this work was corroborated with a smartphone app and web-cum-desktop application for capturing and viewing respectively.
This architecture is based on the idea that prescriptive sampling is vital for useful IBR-algorithms. Hence, it extends the prior plenoptic sampling theory to decrease the dense sampling requirements for traditional light field rendering by decomposing a scene into $D$ depth ranges and individual-range-based light field samples. This allows the camera sampling interval to be increased by a factor of $D$ and the light field spectrum emitted by a scene within each depth range to lie within a densely packed shape that is $D$ times tighter that the full scene’s double-wedge spectrum as depicted in fig. 3. This synthesis pipeline also leaves a possibility for it to be used in the future light field hardwares to reduce required cameras. As discussed by Chai et al [22], the Light Field’s Fourier support (without any occlusions and Lambertian effects [254]) lies inside double-wedge shape within scene depths limit $z_{min},z_{max}]$ as given by the figure 3, whereas occlusion expands the limits of this shape in the light field as the occluder acts as convolver due to the kernel (lying on the line corresponding to the occluder’s depths) with farther scene content. Further, Fourier support of the light field gets limited due to the closest occluder which convolves the furthest scene content’s line resulting in a parallelogram-like-shape as depicted in fig 3. This can only be half packed when compared to double-wedge shape and hence required maximum camera sampling interval $\Delta_{u}$ for an occluded light field and is given be eqn. 57
@@
Figure 3: Traditional Plenoptic Sampling without occlusions [22] in (a) where the double-wedge (in blue) encloses the fourier support of a light field without occlusions. Here, the double-wedge width is used to calculate Nyquist Rate Sampling, that determines by minimum and maximum scene depths $z_{min},z_{max}]$ and $K_{x}$ (max spatial frequency). In (b), the light field is split into $D$ non-overlapping layers and Nyquist Sampling Rate decreases by a factor of $D$ and (c) represents that without occlussions, the full light field spectrum is sum of the spectra from each layer. In LLFF [147] extends this to mitigate occlusions as continuous light fields are reconstructed from MPIs. In (d), occlusions expand the Fourier supports to ——gm (where blue represents support without occlusions and purple regions depict the occlusion expanding the fourier support) doubling the Nyquist Sampling Rate. Further, (e) depicts that separate reconstruction of the light field for $D$ layers decrease the Nyquist Sampling Rate by a factor of $D$ whereas (f) depicts the case that full light spectrum cannot be reconstructed by summation over individual layer spectra because their support union in smaller than support union in (a). Hence, the need for alpha compositing individual light layers from back to front in the primal domain to compute the full light field. Figures used from LLFF original paper [147]
@@
where $f$ is the focal length and $z_{min}$ is the minimum distance to the scene. This strategy for MPI neighborhood blending is particularly efficient for non-Lambertian renderings. Virtual apparent depth of a specularity changes with viewpoint [207] for curved surfaces and hence, specularities can appear as curves in light fields’ epipolar slices and lines can appear for diffuse points. Here, each of the MPI represents a specularity for a local range of views by placing the specularity at a single virtual depth. This approach leverages real world and synthetic dataset of natural scenes [200, 175, 186, 187]. These results that quantitatively validate and demonstrate that LLFFs match the perceptual quality of Nyquist rate Sampling while leveraging $\approx 4000\times$ fewer input images has been demonstrated for over 60 diverse real world scenes and this work paved way for future IBR algorithms.
@@
$\theta:\chi\times\mathbb{R}^{3\times 4}\times\mathbb{R}^{3\times 3}\rightarrow\mathbb{R}^{H\times W\times 3},(\phi,\textrm{{E}},\textrm{{K}})\mapsto\theta(\phi,\textrm{{E}},\textrm{{K}})=\textit{I}$ (67)
@@
[…]
@@
where $\textrm{{V}}_{\textrm{rgb}}(\textbf{{x}})\in\mathbb{R}^{3}$ and $\textrm{{V}}_{\alpha}(\textbf{{x}})$ are the color and opacity (range $0,\infty]$) at the 3D coordinate x. This enables the model to learn soft discrete volume representation that allowed gradients to flow backward easily and model fine and transparent structures (like hair and smoke).
NV also discusses the potential of leveraging only 3 images from 3 orthogonally placed cameras to achieve maximize convergence while downsampling an image by a factor of 8. Here, the Encoder is a Convolutional Variational encoder whereas the authors leverage an MLP for Decoder (which demands prohibitive memory requirement to produce reconstructions with higher quality). Hence, Lombardi et al [135] model the volume function as a discrete 3D grid of voxels hence, gaining the output tensor in the form of discretized voxels. The authors here optimized the decoder as a function $S(\textrm{{x}; {Y}}):\mathbb{R}^{3}\rightarrow\mathbb{R}^{C}$ to samples from the grid Y by scaling continuous values in the range [1,1][-1,1] to $[1,D]$ in each dimension followed by trilinear interpolation. Hence, for general case of 3D cube with center at $\textrm{{x}}_{0}$ with $W$ sides is given by the eqn. 71.
$\textrm{{V(x; z)}}=S\cdot\left(\frac{\textrm{x - x}_{0}}{W/2};\textrm{{g(z)}}\right),$ (71)
@@
$\textrm{where }\alpha(t)=min\left(\int_{t_{min}}^{t}\textrm{{V}}_{\alpha}(\textrm{{r}}_{o}+s\textrm{{r}}_{d})\textrm{ d}s,1\right),$ (77)
[…]
@@
$\textrm{and }\mathcal{P}_{\beta}(\textrm{{I}}_{\alpha})=\frac{1}{P}\sum_{\textrm{{p}}}\lambda_{\beta}\left[\textrm{log({I}${}_{\alpha}$({p})) + log(1 -- {I}${}_{\alpha}$({p}))}\right]$ (81)
@@
[…]
@@
where $[a]$ is nearest integer for $a\in\mathbb{R}$. Further, this approach proposes to leverage a z-buffer to remove occluded points as several points can project onto the same pixel but the lack of topological information in point clouds can result in the bleeding problem (where occluded surfaces and background points could be seen through from front surface). This approach proposes to leverage multi-scale progressive rendering to avoid relying on choice of disk radius (important for splatting which uses a 3D elliptic disk with an estimated radius to replace each point). Here, a sequence of images S[1],S[2],,S[T]S[1],S[2],...,S[T], where the i-th image has the size of $\frac{W}{2^{t}}\times\frac{H}{2^{t}}$ is obtained by rendering a point cloud on a canvas pyramid of different resolution $T$ times by performing simple point-cloud projection. Hence, $S[1]$ has the maximum details and is prone to the most surface bleeding whereas $S[T]$ suffers from least surface bleeding but has the minimum geometric details. Every other image in sequence (S[2],S[3],,S[T1]S[2],S[3],...,S[T-1]) has different detail-bleeding tradeoffs. Now, all the raw sequence of images are mapped to three channel RGB image $I$ using a U-Net [183] (with Gated Convolutions [252] to handle sparse inputs) based rendering network ($\mathcal{R}_{\theta}$ as given in the eqn, 85 below.
@@
The, the raw images $S[i]$ are concatenated to the U-Net first block with corresponding resolutions (this coarse-to-fine mechanism and mipmapping  are similar). Hence, rendering network provides the mechanism for implicit detail selection in this case. Further, OpenGL is used for rasterization of raw image sequences S[1],S[2],,S[T]S[1],S[2],...,S[T]. The dimensionality of the descriptors is set to eight ($M=8$).

### v2

@@ Abstract
Keywords:
NeRFs (Neural Radiance Fields) Rendering Computer Vision Deep Learning Computer Graphics Robotics Augmented Reality Virtual Reality Implicit Machine Learning
@@ 1 Introduction
The §1 presents a brief introduction that this paper aims to resolve i.e. to delineate a classification for NeRF models and discuss about its history and future. The remainder of this paper has been organized as follow. §3 discusses the various neural architectures before the conceptions of NeRFs. It further discusses various surveys that have been put forward in this domain and the need for this survey. §4 discusses the chronological description of the models, extensions, and improvements over Neural Radiance Fields. §5 briefly discusses the datasets that have been leveraged till now in the field of NeRFs. Further, these are also then taxonomized according to different objectives (discussed in §6), different evaluation metrics discussed in §7), and different applications (discussed in §8). All sections from §58 discuss the research in the past couple of years in the field of NeRFs (limited till December 2022). Finally, the future prospects of this family of models is explored in §9. §10 concludes the paper and discusses some future prospects for this research.
@@ 2.1 Computer Graphics
    $\mathcal{N}=\frac{\Delta f(x,y,z)}{||\Delta f(x,y,z)||}$ (3)
    where $||v||$ denotes the magnitude of vector $v$. The resulting $\mathcal{N}$ vector is a unit vector that points in the direction of the surface normal.
@@
      $\textrm{H}=\begin{cases}0\degree,\Delta=0,\\ 60\degree\times\left(\frac{\textrm{G}^{\prime}-\textrm{B}^{\prime}}{\Delta}mod6\right),\textrm{C}_{max}=\textrm{R}^{\prime}\\ 60\degree\times\left(\frac{\textrm{B}^{\prime}-\textrm{R}^{\prime}}{\Delta}\textrm{ }+2\right),\textrm{ }\textrm{C}_{max}=\textrm{G}^{\prime}\\ 60\degree\times\left(\frac{\textrm{R}^{\prime}-\textrm{G}^{\prime}}{\Delta}\textrm{ }+4\right),\textrm{ }\textrm{C}_{max}=\textrm{B}^{\prime}\\ \end{cases}$ (21)
      $\textrm{S}=\begin{cases}0\degree,\textrm{ }\textrm{C}_{max}=0,\\ \frac{\Delta}{\textrm{C}_{max}},\textrm{C}_{max}\not=0\\ \end{cases}$ (22)
@@
    However, when the pitch angle approaches ±$\pm 90\degree$, the rotation around the y-axis (yaw) and the rotation around the z-axis (roll) become aligned, resulting in a loss of a degree of freedom. This is gimbal lock, and it can cause unexpected or incorrect behavior in animations and simulations.
@@ 2.2 Mathematics and Geometry
      $\mathcal{NSDF}(p)=\frac{f(p)}{||f(p)||_{2}}$ (37)
      where $p$ is 3D point, $\textrm{{MLP}}(p)$ is MLP output applied to $p$, and $||f(p)||_{2}$ is the L2 norm of the MLP output. The NSDF output is a unit vector that points towards the nearest surface of the object. There are different types of NSDF used in computer vision and robotics applications. Some of them are:
@@
    $\mathcal{G}(x)=e^{-\frac{||x||^{2}}{\sigma^{2}}}$ (41)
    where $x$ is 3D vector for point in space, $||x||$ is its Euclidean distance, and $\sigma$ scales and controls the size of Gaussiann function.
@@
  • Discrete Fourier Transform is a mathematical technique that is a discrete variant of the Fourier Transform discussed above. It transforms a discrete-time sequence of N complex numbers $x[n]$ into a sequence of N complex numbers $X[k]$ representing the frequency content of the signal. It is defined by the eqn 47 where $k$ is frequency index ranging from 00 to $N-1$.
@@
    It can be calculated using matrix multiplication of the $N\times N$ DFT matrix $F$ and the $N-1$ column vectors.
@@
    The equation for positional encoding is given in eqn 48 where $\text{PE}{(pos,2i)}$ and $\text{PE}{(pos,2i+1)}$ are the $2i$-th and $(2i+1)$-th elements of the positional encoding vector for the $pos$-th position in the sequence, $d_{\text{model}}$ is the dimensionality of the model, and $i$ is the index of the embedding dimension.
@@ 3 History of Neural Rendering and NeRFs
LLFF (Local Light Field Fusion) (example depicted in fig. 3(a)) is one prior work done by Mildenhall et al [148], some of whose authors were also contributors for the work under focus in this survey. LLFFs leverage MPI (Multiplane Image) scene representations [274] for view synthesis from an irregular grid of sampled views for expanding each sampled view into a local light fields [122] and then renders novel views by fusing these fields. LLFFs achieve perceptual quality similar to Nyquist Rate Sampling (with 4000×4000$\times$4000 × less views) by creating a bound for real world scenes creation/rendering which is specifies the density of sample views for a given scene by extending the traditional plenoptic sampling [22] theory. Here, the LLFFs result in a quadratic decrease in the number of input views based on the number of predicted planes for each layered scene representation and still maintain the Nyquist level performance by specializing to the subset of natural scenes. Further, this work was corroborated with a smartphone app and web-cum-desktop application for capturing and viewing respectively.
This architecture is based on the idea that prescriptive sampling is vital for useful IBR-algorithms. Hence, it extends the prior plenoptic sampling theory to decrease the dense sampling requirements for traditional light field rendering by decomposing a scene into $D$ depth ranges and individual-range-based light field samples. This allows the camera sampling interval to be increased by a factor of $D$ and the light field spectrum emitted by a scene within each depth range to lie within a densely packed shape that is $D$ times tighter that the full scene’s double-wedge spectrum as depicted in fig. 3. This synthesis pipeline also leaves a possibility for it to be used in the future light field hardwares to reduce required cameras. As discussed by Chai et al [22], the Light Field’s Fourier support (without any occlusions and Lambertian effects [258]) lies inside double-wedge shape within scene depths limit $[z_{min},z_{max}]$ as given by the figure 3, whereas occlusion expands the limits of this shape in the light field as the occluder acts as convolver due to the kernel (lying on the line corresponding to the occluder’s depths) with farther scene content. Further, Fourier support of the light field gets limited due to the closest occluder which convolves the furthest scene content’s line resulting in a parallelogram-like-shape as depicted in fig 3. This can only be half packed when compared to double-wedge shape and hence required maximum camera sampling interval $\Delta_{u}$ for an occluded light field and is given be eqn. 57
@@
Figure 3: Traditional Plenoptic Sampling without occlusions [22] in (a) where the double-wedge (in blue) encloses the fourier support of a light field without occlusions. Here, the double-wedge width is used to calculate Nyquist Rate Sampling, that determines by minimum and maximum scene depths $[z_{min},z_{max}]$ and $K_{x}$ (max spatial frequency). In (b), the light field is split into $D$ non-overlapping layers and Nyquist Sampling Rate decreases by a factor of $D$ and (c) represents that without occlussions, the full light field spectrum is sum of the spectra from each layer. In LLFF [148] extends this to mitigate occlusions as continuous light fields are reconstructed from MPIs. In (d), occlusions expand the Fourier supports to ——gm (where blue represents support without occlusions and purple regions depict the occlusion expanding the fourier support) doubling the Nyquist Sampling Rate. Further, (e) depicts that separate reconstruction of the light field for $D$ layers decrease the Nyquist Sampling Rate by a factor of $D$ whereas (f) depicts the case that full light spectrum cannot be reconstructed by summation over individual layer spectra because their support union in smaller than support union in (a). Hence, the need for alpha compositing individual light layers from back to front in the primal domain to compute the full light field. Figures used from LLFF original paper [148]
@@
where $f$ is the focal length and $z_{min}$ is the minimum distance to the scene. This strategy for MPI neighborhood blending is particularly efficient for non-Lambertian renderings. Virtual apparent depth of a specularity changes with viewpoint [209] for curved surfaces and hence, specularities can appear as curves in light fields’ epipolar slices and lines can appear for diffuse points. Here, each of the MPI represents a specularity for a local range of views by placing the specularity at a single virtual depth. This approach leverages real world and synthetic dataset of natural scenes [202, 177, 188, 189]. These results that quantitatively validate and demonstrate that LLFFs match the perceptual quality of Nyquist rate Sampling while leveraging 4000$\approx$ 4000$\times$ 4000 × fewer input images has been demonstrated for over 60 diverse real world scenes and this work paved way for future IBR algorithms.
@@
[…]
@@
$\end{aligned}\end{multlined}\operatorname*{arg\,min}_{\{\theta,\Psi,\{\textrm{{z}}_{j}\}^{m}_{j=1}\}}\sum_{j=1}^{M}\sum_{i=1}^{N}\underbrace{||\Theta_{\theta}(\Phi_{\psi(\textrm{{z}}_{j})},\textbf{E}_{i}^{j},\textbf{K}_{i}^{j})||_{2}^{2}}_{\mathcal{L}_{\textrm{img}}}\\ \begin{aligned} &+\underbrace{\lambda_{depth}||\textrm{min({d}}_{i,final}^{j},0)||_{2}^{2}}_{\mathcal{L}_{\textrm{depth}}}\\ &+\underbrace{\lambda_{latent}||\textrm{{z}}_{j}||_{2}^{2}}_{\mathcal{L}_{\textrm{latent}}}\end{aligned}$ (69)
@@
where $\textrm{{V}}_{\textrm{rgb}}(\textbf{{x}})\in\mathbb{R}^{3}$ and $\textrm{{V}}_{\alpha}(\textbf{{x}})$ are the color and opacity (range $[0,\infty]$) at the 3D coordinate x. This enables the model to learn soft discrete volume representation that allowed gradients to flow backward easily and model fine and transparent structures (like hair and smoke).
NV also discusses the potential of leveraging only 3 images from 3 orthogonally placed cameras to achieve maximize convergence while downsampling an image by a factor of 8. Here, the Encoder is a Convolutional Variational encoder whereas the authors leverage an MLP for Decoder (which demands prohibitive memory requirement to produce reconstructions with higher quality). Hence, Lombardi et al [136] model the volume function as a discrete 3D grid of voxels hence, gaining the output tensor in the form of discretized voxels. The authors here optimized the decoder as a function $S(\textrm{{x}; {Y}}):\mathbb{R}^{3}\rightarrow\mathbb{R}^{C}$ to samples from the grid Y by scaling continuous values in the range $[-1,1]$ to $[1,D]$ in each dimension followed by trilinear interpolation. Hence, for general case of 3D cube with center at $\textrm{{x}}_{0}$ with $W$ sides is given by the eqn. 71.
[…]
@@
where $\textrm{where }\alpha(t)=min\left(\int_{t_{min}}^{t}\textrm{{V}}_{\alpha}(\textrm{{r}}_{o}+s\textrm{{r}}_{d})\textrm{ d}s,1\right),$ (77)
$\textrm{and }\textrm{{r}}_{d}=\frac{\textrm{{P}}^{-1}\textrm{{p--r}}_{o}}{||\textrm{{P}}^{-1}\textrm{{p--r}}_{o}||}\in\mathbb{R}^{3},$ (78)
@@
and $\textrm{and }\mathcal{P}_{\beta}(\textrm{{I}}_{\alpha})=\frac{1}{P}\sum_{\textrm{{p}}}\lambda_{\beta}\left[\textrm{log({I}${}_{\alpha}$({p})) + log(1 -- {I}${}_{\alpha}$({p}))}\right]$) ] (81)
@@
$\end{aligned}\end{multlined}l(\theta)=\frac{1}{P}\sum_{\textrm{{p}}}\left|\left|\widehat{\textrm{{I}}}_{\textrm{rgb}}(\textrm{{p}})\textbf{--}\textrm{{I}}_{\textrm{rgb}}^{*}(\textrm{{p}})\right|\right|^{2}\\ \begin{aligned} &+\lambda_{\textrm{KL}}D_{\textrm{KL}}\left(\textrm{{z}}||\mathcal{N}(0,1)\right)\\ &+\mathcal{P}_{\textrm{var\_lvoxop}}(\textrm{{V}}_{\alpha})\\ &+\mathcal{P}_{\beta}(\textrm{{I}}_{\alpha}).\end{aligned}$ (82)
@@
where $[a]$ is nearest integer for $a\in\mathbb{R}$. Further, this approach proposes to leverage a z-buffer to remove occluded points as several points can project onto the same pixel but the lack of topological information in point clouds can result in the bleeding problem (where occluded surfaces and background points could be seen through from front surface). This approach proposes to leverage multi-scale progressive rendering to avoid relying on choice of disk radius (important for splatting which uses a 3D elliptic disk with an estimated radius to replace each point). Here, a sequence of images $S[1],S[2],...,S[T]$, where the i-th image has the size of $\frac{W}{2^{t}}\times\frac{H}{2^{t}}$ is obtained by rendering a point cloud on a canvas pyramid of different resolution $T$ times by performing simple point-cloud projection. Hence, $S[1]$ has the maximum details and is prone to the most surface bleeding whereas $S[T]$ suffers from least surface bleeding but has the minimum geometric details. Every other image in sequence ($S[2],S[3],...,S[T-1]$) has different detail-bleeding tradeoffs. Now, all the raw sequence of images are mapped to three channel RGB image $I$ using a U-Net [185] (with Gated Convolutions [255] to handle sparse inputs) based rendering network ($\mathcal{R}_{\theta}$ as given in the eqn, 85 below.
@@
The, the raw images $S[i]$ are concatenated to the U-Net first block with corresponding resolutions (this coarse-to-fine mechanism and mipmapping are similar). Hence, rendering network provides the mechanism for implicit detail selection in this case. Further, OpenGL is used for rasterization of raw image sequences $S[1],S[2],...,S[T]$. The dimensionality of the descriptors is set to eight ($M=8$).